A cannon from the Napoleonic wars was able to launch cannon balls at a velocity of 540 m s–1. The cannon is fired at the top of a cliff with the barrel in the horizontal position.
(a) Neglecting air resistance, what is the range of the cannon ball, to the point where it hits the water 100 m below?
(4 marks)
(b) Calculate the initial horizontal component of the momentum of the cannon ball, if it has a mass of 5.44 kg. (2 marks)
(c) If the cannon has a mass of 5000 kg and is on a trolley on the horizontal surface, what is its initial recoil velocity?
(2 marks)
1
Expert's answer
2020-12-23T07:35:19-0500

(a)Time of falling "t" can be found from the kinematic equation (if there is no vertical initial velocity component):



"h = \dfrac{gt^2}{2}"

where "h = 100m" is the height, and "g = 9.8m\/s^2" is the gravitational acceleration. Thus, obtain:


"t = \sqrt{\dfrac{2h}{g}} = \sqrt{\dfrac{2\cdot 100}{9.8}} \approx 4.5s"

In this time the stone covers the following horizontal distance:


"d = v_0t = 540m/s\cdot 4.5s\approx 2439m"

Here "v_0 = 540m\/s" is the initial horizontal velocity of the ball.


(b) The initial horizontal component of the momentum of the cannon ball of mass "m = 5.44kg" is given by the expression:


"p_{ball} = mv_0\\
p_{ball} = 5.44\cdot 540 = 2937.6\space kg\cdot m/s"



(c) The initial recoil momentum is equal to the initial horizontal component of the momentum of the cannon ball (according to the momentum conservation):


"p_{recoil} = p_{ball}\\
m_{recoil}v_{recoil} = p_{ball}\\"

where "m_{recoil} = 5000kg" is the mass of the cannon. Then the velocity of the recoil is:


"v_{recoil} = \dfrac{p_{ball}}{m_{recoil}}\\
v_{recoil} = \dfrac{2937.6}{5000} \approx 0.59m/s"

Answer. (a) 2439 m, (b) 2937.6 kg*m/s, (c) 0.59 m/s.